A proton moving at v1 = 7.9Mm/s collides elastically and head-on with a second proton moving in the opposite direction at v2 = 8.6Mm/s. Choose positive velocities in the direction of v? 1.

a.

Find the velocity of the first proton after the collision

b.

Find the velocity of the second proton after the collision.
Answer #1

m1=m2 =m=1.67x10^-27 kg

u1=7.9x10^6 m/s

u2= -8.6x10^6 m/s

after the collision

v1=(m1-m2)*u1/(m1+m2) +2m2u2/(m1+m2)=2mx-8.6x10^6/(2m)=-8.6x106 m/s

second proton

v2 =(m2-m1)*u2/(m1+m2)+2m1u1/(m1+m2) =2m*7.9x10^6/(2m) =7.9x10^6 m/s